OpenStudy (anonymous):

Two jars of different brand of strawberry jam stand on a shelf in a supermarket. the taller jar is twice the height of the other jar, but its diameter is one half as much as the diameter of the shorter jar. if the shorter jar costs twice as much as the taller jar ehich is the better buy, assuming equal quality?

OpenStudy (anonymous):

The other jar. I will tell why. \[ T=\frac{1}{4} \pi d^2 h\\ O=2 \pi d^2 h \] Can you conclude?

OpenStudy (anonymous):

@christine.joy what's the volume of the first jar and that of the second jar, compare the two and you'll get the better buy ;)

OpenStudy (anonymous):

\[ \frac OT =8 \]

OpenStudy (anonymous):

volume of a cylinder = pi * r^2 * h

OpenStudy (anonymous):

you can make it easier on yourself @christine.joy if you assumed the taller jar is 20 cm high with a diameter of 4 cm, then the shorter jar is 10cm high with a diameter of 8cm, now using the above volume formula, which jar has more jam in it ?

OpenStudy (anonymous):

based on the above comment I get v1 = 80pi volume of the taller jar v2 = 160pi volume of the shorter jar

OpenStudy (anonymous):

so really both are the same value when buying, the shorter jar costs double the taller jar but at the same time it contains double the jam quantity